Tohban Report 2012-03-14
From RHESSI Wiki
|Start Date:||7 Feb 2012|
|End Date:||14 Feb 2012|
|Tohban:||Juan Carlos Martinez Oliveros|
|Next Tohban:||Hazel Bain|
|List all reports|
Solar activity was moderate. There are currently 6 sunspots on disk. No Solar activity is expected to continue at the current rate in the next week.
How many GOES flares occurred?
Flares above B, C, M, X class were 3 36 3 2
And how many of these are listed in the RHESSI flare list?
Flares above B, C, M, X class were 0 22 3 0
And how many had EXCELLENT coverage?
Flares above B, C, M, X class were 0 10 0 0
There were RHESSI flares/GOES flares 210 / 44 over the time range 07-Mar-12 14-Mar-12
During the week the SSR steadily dropped, as was predicted. By March 9 the SSR was emptying at end of each pass. During the last days of the week two M class flares occurred filling the SSR, reaching a maximum of ~35%.
During this week operations were dominated by the activation of functions disabled during the high SSR incident last week.
The thick attenuators were enabled for automatic motion on Wednesday 7 at 20:20:32 UTC. Also on this day the thresholds for the G2 and G4 detectors were restore to the normal values. At this point the thin attenuator was lock in and detectors G2 and G4 were not taking events during HLA crossings.
The steps mentioned above went smoothly and with the massive number of passes the SSR level dropped steadily. By March 8 the SSR was less than 5%.
On March 9 at 01:19:49 UTC the thin shutter were enable to move, returning RHESSI to its nominal attenuator allowed states of 0-1-3.
On Monday, March 12 at 16:33:20 UTC, the commanding to turn off G2 and G4 events during high latitude crossings were removed from active ATS. The SSR has been emptying after the Berkeley pass sets even after the large M8.4 flare on the 10th. Also we are back to the normal ~10 passes per day.
On the Wednesday ops meeting was decided to change the decimation to normal/vigorous. This is the last change to undo after the, shutter/fast rate/high SSR fill level, anomaly.
Some noise was observed on detector 9 during the week. On March 8 at 01:24 UTC detector G9 HV was lowered by ten counts. Some sporadic noise was observed, but the detector appears to fixed it-self.
GAPS IN APP_ID = 102 (VC3-MONITOR RATES) WITH PACKET RATE LT 30 N_GAPS 1 GAP START TIME GAP END TIME GAP (SEC) 2012-03-07T04:50:00.000 -- 2012-03-07T06:40:00.000 6600.0000
DATA GAPS FOR 2012/03/12 TIME RANGE: 2012-03-12T00:00:00.000 -- 2012-03-13T00:00:00.000 GAPS IN APP_ID = 1 (VC1-SOH) WITH PACKET RATE LT 30 N_GAPS 1 GAP START TIME GAP END TIME GAP (SEC) 2012-03-12T02:50:00.000 -- 2012-03-12T02:55:00.000 300.00000
(VC3-MONITOR RATES) WITH PACKET RATE LT 30 N_GAPS 1 GAP START TIME GAP END TIME GAP (SEC) 2012-03-12T16:30:00.000 -- 2012-03-12T16:35:00.000 300.00000
Another instance when the SAS failed to provide a Sun vector was observed on March 7 at ~00:30 UTC. This behavior was observed only one orbit and later the SAS corrected itself.